# A Survey on 3D Gaussian Splatting

### v7

@@ 6 Performance Comparison
TABLE III: Comparison of mapping methods (§6.2) on Replica [261] (static scenes), in terms of PSNR, SSIM, and LPIPS. The results for FPS are taken from [113] using one 4090 GPU.
@@
In this section, we provide more empirical evidence by presenting the performance of several 3D GS algorithms that we previously discussed. The diverse applications of 3D GS across numerous tasks, coupled with the custom-tailored algorithmic designs for each task, render a uniform comparison of all 3D GS algorithms across a single task or dataset impracticable. For comprehensiveness, we provide a collection of representative datasets in Table 6 according to our analysis in Sec. 5. Due to the limited space, we have chosen several representative tasks for an in-depth performance evaluation. The performance scores are primarily sourced from the original papers, except where indicated otherwise. We also maintain a Github repository for this section.
6.1 Performance Benchmarking: Localization
$\bullet$ Dataset: Replica [261] dataset is a collection of 18 highly detailed 3D indoor scenes. These scenes are not only visually realistic but also offer comprehensive data including dense meshes, high-quality HDR textures, and detailed semantic information for each element. Following [262], three sequences about rooms and five sequences about offices are used for the evaluation.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent 3D GS based algorithms [114, 113, 112, 111] and six typical SLAM methods [262, 263, 264, 265, 266, 267].
$\bullet$ Evaluation Metric: The root mean square error (RMSE) of the absolute trajectory error (ATE) is a commonly used metric in evaluating SLAM systems [275], which measures the root mean square of the Euclidean distances between the estimated and true positions over the entire trajectory.
$\bullet$ Result: As shown in Table 5.6, the recent 3D Gaussians based localization algorithms have a clear advantage over existing NeRF based dense visual SLAM. For example, SplaTAM [112] achieves a trajectory error improvement of $\sim$50%, decreasing it from 0.52cm to 0.36cm compared to the previous state-of-the-art (SOTA) [266]. We attribute this to the dense and accurate 3D Gaussians reconstructed for scenes, which can handle the noise of real sensors. This reveals that effective scene representations can improve the accuracy of localization tasks.
6.2 Performance Benchmarking: Static Scenes
@@
$\bullet$ Dataset: The same dataset as in Sec. 6.1, i.e., Replica [261], is used for comparison. The testing views are the same as those collected by [262].
@@
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent papers which introduce 3D Gaussians into their systems [114, 113, 112, 111], as well as three dense SLAM methods [264, 263, 266].
@@
$\bullet$ Evaluation Metric: Peak signal-to-noise ratio (PSNR), structural similarity (SSIM) [300], and learned perceptual image patch similarity (LPIPS) [301] are used for measuring RGB rendering performance.
@@
$\bullet$ Result: Table 6 shows that 3D Gaussians based systems generally outperform the three dense SLAM competitors. For example, Gaussian-SLAM [114] establishes new SOTA and outperforms previous methods by a large margin. Compared to Point-SLAM [266], GSSLAM [113] is about 578 times faster in achieving very competitive accuracy. In contrast to previous method [266] that relies on depth information, such as depth-guided ray sampling, for synthesizing novel views, 3D GS based system eliminates this need, allowing for high-fidelity rendering for any views.
6.3 Performance Benchmarking: Dynamic Scenes
@@
$\bullet$ Dataset: D-NeRF [184] dataset includes videos with 50 to 200 frames each, captured from unique viewpoints. It features synthetic, animated objects in complex scenes, with non-Lambertian materials. The dataset provides 50 to 200 training images and 20 test images per scene, designed for evaluating models in the monocular setting. The testing views are the same as the original paper [184].
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve five recent papers that model dynamic scenes with 3D GS [126, 95, 132, 94, 93], as well as six NeRF based approaches [184, 302, 37, 303, 187, 304].
$\bullet$ Evaluation Metric: The same metrics as in Sec. 6.2, i.e., PSNR, SSIM [300], and LPIPS [301], are used for evaluation.
$\bullet$ Result: From Table 6.3 we can observe that 3D GS based methods outperform existing SOTAs by a clear margin. The static version of 3D GS [10] fails to reconstruct dynamic scenes, resulting in a sharp drop in performance. By modeling the dynamics, D-3DGS [94] outperforms the SOTA method, FFDNeRF [187], by 6.83dB in terms of PSNR. These results indicate the effectiveness of introducing additional properties or structured information to model the deformation of Gaussians so as to model the scene dynamics.
TABLE IV: Comparison of reconstruction methods (§6.3) on D-NeRF [184] (dynamic scenes), in terms of PSNR, SSIM, and LPIPS. denotes results reported in [95].
@@
D-NeRF [184] [CVPR21] 30.50 0.95 0.07
TiNeuVox-B [302] [SGA22] 32.67 0.97 0.04
KPlanes [37] [CVPR23] 31.61 0.97 -
HexPlane-Slim [303] [CVPR23] 32.68 0.97 0.02
FFDNeRF [187] [ICCV23] 32.68 0.97 0.02
MSTH [304] [NeurIPS23] 31.34 0.98 0.02
3D GS [10] [TOG23] 23.19 0.93 0.08
4DGS [93] [ICLR24] 34.09 0.98 -
4D-GS [95] [CVPR24] 34.05 0.98 0.02
GaGS [132] [CVPR24] 37.36 0.99 0.01
CoGS [126] [CVPR24] 37.90 0.98 0.02
D-3DGS [94] [CVPR24] 39.51 0.99 0.01
6.4 Performance Benchmarking: Human Avatar
$\bullet$ Dataset: ZJU-MoCap [292] is a prevalent benchmark in human modeling from videos, captured with 23 synchronized cameras at a $\times1024$ resolution. Six subjects (i.e., 377, 386, 387, 392, 393, and 394) are used for evaluation [305]. The same testing views following [306] are adopted.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve three recent papers which model human avatar with 3D GS [146, 145, 249], as well as six human rendering approaches [292, 307, 308, 309, 305, 306].
$\bullet$ Evaluation Metric: PSNR, SSIM [300], and LPIPS* [301] are used for measuring RGB rendering performance. Here LPIPS* equals to LPIPS $\times$ 1000.
$\bullet$ Result: Table 6.5 presents the numerical results of top-leading solutions in human avatar modeling. We observe that introducing 3D GS into the framework leads to consistent performance improvements in both rendering quality and speed. For instance, GART [146] outperforms current SOTA, Instant-NVR [306], by 1.21dB in terms of PSNR. Considering the enhanced fidelity, inference speed and editability, 3D GS based avatar modeling may revolutionize the field of 3D animation, interactive gaming, etc.
6.5 Performance Benchmarking: Surgical Scenes
3D reconstruction from endoscopic video is critical to robotic-assisted minimally invasive surgery, enabling preoperative planning, training through AR/VR simulations, and intraoperative guidance.
$\bullet$ Dataset: EndoNeRF [258] dataset presents a specialized collection of stereo camera captures, comprising two samples of in-vivo prostatectomy. It is tailored to represent real-world surgical complexities, including challenging scenes with tool occlusion and pronounced non-rigid deformation. The same testing views as in [260] are used.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve three recent papers which reconstruct dynamic 3D endoscopic scenes with GS [152, 153, 155], as well as three NeRF-based surgical reconstruction approaches [258, 259, 260].
$\bullet$ Evaluation Metric: PSNR, SSIM [300], and LPIPS [301] are adopted for evaluation. In addition, the requirement for GPU memory is also reported.
TABLE V: Comparison of reconstruction methods (§6.4) on ZJU-MoCap [292] (avatar), in terms of PSNR, SSIM, and LPIPS*. The results for non-GS methods are taken from [146].
[…]
NeuralBody [292] [CVPR21] 29.03 0.96 42.47
AnimNeRF [307] [ICCV21] 29.77 0.96 46.89
PixelNeRF [308] [ICCV21] 24.71 0.89 121.86
NHP [309] [NeurIPS21] 28.25 0.95 64.77
HumanNeRF [305] [CVPR22] 30.66 0.97 33.38
Instant-NVR [306] [CVPR23] 31.01 0.97 38.45
GauHuman [145] [CVPR24] 31.34 0.97 30.51
3DGS-Avatar [249] [CVPR24] 30.61 0.97 29.58
GART [146] [CVPR24] 32.22 0.98 29.21
TABLE VI: Comparison of reconstruction methods (§6.5) on EndoNeRF [258] (surgical scenes), in terms of PSNR, SSIM, and LPIPS. The results for non-GS methods are taken from [153]. FPS and GPU usage for training (Mem.) are measured using one 4090 GPU [153].
[…]
EndoNeRF [258] [MICCAI22] 36.06 0.93 0.09 0.04 19GB
EndoSurf [260] [MICCAI23] 36.53 0.95 0.07 0.04 17GB
LerPlane-9k [259] [MICCAI23] 35.00 0.93 0.08 0.91 20GB
LerPlane-32k [259] [MICCAI23] 37.38 0.95 0.05 0.87 20GB
Endo-4DGS [152] [MICCAI24] 37.00 0.96 0.05 - 4GB
EndoGaussian [153] [arXiv] 37.85 0.96 0.05 195.09 2GB
HFGS [155] [BMVC24] 38.14 0.97 0.03 - -
$\bullet$ Result: Table 6.5 shows that introducing the explicit representation of 3D Gaussians leads to several significant improvements. For instance, EndoGaussian [153] outperforms a strong baseline, LerPlane-32k [259], among all metrics. In particular, EndoGaussian demonstrates an approximate 224-fold increase in speed while consumes just 10% of the GPU resources. These impressive results attest to the efficiency of GS-based methods, which not only expedite processing but also minimize GPU load, thus easing the demands on hardware. Such attributes are vitally significant for real-world surgical application deployment, where optimized resource usage can be a key determinant of practical utility.
7 Future Research Directions
@@
$\bullet$ Physics- and Semantics-aware Scene Representation. As a new, explicit scene representation technique, 3D Gaussian offers transformative potential beyond merely enhancing novel-view synthesis. It has the potential to pave the way for simultaneous advancements in scene reconstruction and understanding by devising physics- and semantics-aware 3D GS systems. While significant progress has been made in physics (Sec. 5.7) and semantics [310, 311, 312, 313, 314, 315] individually, there remains considerable untapped potential in their synergistic integration. This is poised to revolutionize a range of fields and downstream applications. For instance, incorporating prior knowledge such as the general shape of objects can reduce the need for extensive training viewpoints [48, 47] while improving geometry/surface reconstruction [316, 77]. A critical metric for assessing scene representation is the quality of its generated scenes, which encompasses challenges in geometry, texture, and lighting fidelity [128, 141, 66]. By merging physical principles and semantic information within the 3D GS framework, one can expect that the quality will be enhanced, thereby facilitating dynamics modeling [166, 21], editing [92, 90], generation [133, 134], and beyond. In a nutshell, pursuing this advanced and versatile scene representation opens up new possibilities for innovation in computational creativity and practical applications across diverse domains.
@@
$\bullet$ Learning Physical Priors from Large-scale Data. As we explore the potential of physics- and semantics-aware scene representations, leveraging large-scale datasets to learn generalizable, physical priors emerges as a promising direction. The goal is to model the inherent physical properties and dynamics embedded within real-world data, transforming them into actionable insights that can be applied across various domains such as robotics and visual effects. Establishing a learning framework for extracting these generalizable priors enables the application of these insights to specific tasks in a few-shot manner. For instance, it allows for rapid adaptation to new objects and environments with minimal data input. Furthermore, integrating physical priors can enhance not only the accuracy and quality of generated scenes but also their interactive and dynamic qualities. This is particularly valuable in AR/VR environments, where users interact with virtual objects that behave in ways consistent with their real-world counterparts. However, the existing body of work on capturing and distilling physics-based knowledge from extensive 2D and 3D datasets remains sparse. Notable efforts in related area include the continuum mechanics based GS systems (Sec. 5.7), and the generalizable Gaussian representation based on multi-view stereo [317]. Further exploration on real2sim and sim2real might offer viable routes for advancements in this field.
@@
$\bullet$ Modeling Internal Structures of Objects with 3D GS. Despite the ability of 3D GS to produce highly photorealistic renderings, modeling internal structures of objects (e.g., for a scanned object in computed tomography) within the current GS framework presents a notable challenge. Due to the splatting and density control process, the current representation of 3D Gaussian is unorganized and cannot align well with the object’s actual internal structures. Moreover, there is a strong preference in various applications to depict objects as volumes (e.g., computed tomography). However, the disordered nature of 3D GS makes volume modeling particularly difficult. Li et al. [318] used 3D Gaussians with density control as the basis for the volumetric representation and did not involve the splatting process. X-Gaussian [319] involves the splatting process for fast training and inference but cannot generate volumetric representation. Using 3D GS to model the internal structures of objects remains unanswered and deserves further exploration.
@@
$\bullet$ 3D GS for Simulation in Autonomous Driving and beyond. Collecting real-world datasets for autonomous driving is both expensive and logistically challenging, yet crucial for training effective image-based perception systems. To mitigate these issues, simulation emerges as a cost-effective alternative, enabling the generation of synthetic datasets across diverse environments. However, the development of simulators capable of producing photorealistic and diverse synthetic data is fraught with challenges. These include achieving a high level of quality, accommodating various control methods, and accurately simulating a range of lighting conditions. While early efforts [188, 189, 190] in reconstructing urban/street scenes with 3D GS have been encouraging, they are just the tip of the iceberg in terms of the full capabilities. There remain numerous critical aspects to be explored, such as the integration of user-defined object models, the modeling of physics-aware scene changes (e.g., the rotation of vehicle wheels), and the enhancement of controllability and quality (e.g., in varying lighting conditions). Mastery of these capabilities would not only advance autonomous systems but also redefine computational understanding of physical spaces — a leap with implications for world models, spatial intelligence, embodied AI, and beyond.
@@
$\bullet$ Empowering 3D GS with More Possibilities. Despite the significant potential of 3D GS, the full scope of applications for 3D GS remains largely untapped. A promising avenue for exploration involves augmenting 3D Gaussians with additional attributes (e.g., linguistic and spatiotemporal properties as mentioned in Sec. 4.5) and introducing structured information (e.g., spatial MLPs and grids as mentioned in Sec. 4.6), tailored for specific applications. Moreover, recent studies have begun to unveil the capability of 3D GS in several domains, e.g., point cloud registration [320], image representation and compression [60], and fluid synthesis [171]. These findings highlight a significant opportunity for interdisciplinary scholars to explore 3D GS further.
8 Conclusions

### v8

@@ 6.2. Performance Benchmarking: Static Scenes
Table 3. Comparison of mapping methods on Replica (Straub et al., 2019) (static scenes), in terms of PSNR, SSIM, and LPIPS. The results for FPS are taken from (Matsuki et al., 2024) using one 4090 GPU.
@@
$\bullet$ Dataset: The same dataset as in Sec. 6.1, i.e., Replica (Straub et al., 2019), is used for comparison. The testing views are the same as those collected by (Sucar et al., 2021).
@@
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent papers which introduce 3D Gaussians into their systems (Yugay et al., 2023; Matsuki et al., 2024; Keetha et al., 2024; Yan et al., 2024c), as well as three dense SLAM methods (Zhu et al., 2022; Sandström et al., 2023).
@@
$\bullet$ Evaluation Metric: Peak signal-to-noise ratio (PSNR), structural similarity (SSIM) (Wang et al., 2004), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) are used for measuring RGB rendering performance.
@@
$\bullet$ Result: Table 6.2 shows that 3D Gaussians based systems generally outperform the three dense SLAM competitors. For example, Gaussian-SLAM (Yugay et al., 2023) establishes new SOTA and outperforms previous methods by a large margin. Compared to Point-SLAM (Sandström et al., 2023), GSSLAM (Matsuki et al., 2024) is about 578 times faster in achieving very competitive accuracy. In contrast to previous method (Sandström et al., 2023) that relies on depth information, such as depth-guided ray sampling, for synthesizing novel views, 3D GS based system eliminates this need, allowing for high-fidelity rendering for any views.
6.3. Performance Benchmarking: Dynamic Scenes
Table 4. Comparison of reconstruction methods on D-NeRF (Pumarola et al., 2021) (dynamic scenes), in terms of PSNR, SSIM, and LPIPS. denotes results reported in (Wu et al., 2024c).
[…]
D-NeRF (Pumarola et al., 2021) [CVPR21] 30.50 0.95 0.07
TiNeuVox-B (Fang et al., 2022) [SGA22] 32.67 0.97 0.04
KPlanes (Fridovich-Keil et al., 2023) [CVPR23] 31.61 0.97 -
HexPlane-Slim (Cao and Johnson, 2023) [CVPR23] 32.68 0.97 0.02
FFDNeRF (Guo et al., 2023) [ICCV23] 32.68 0.97 0.02
MSTH (Wang et al., 2023a) [NeurIPS23] 31.34 0.98 0.02
3D GS (Kerbl et al., 2023) [TOG23] 23.19 0.93 0.08
4DGS (Yang et al., 2024d) [ICLR24] 34.09 0.98 -
4D-GS (Wu et al., 2024c) [CVPR24] 34.05 0.98 0.02
GaGS (Lu et al., 2024a) [CVPR24] 37.36 0.99 0.01
CoGS (Yu et al., 2024b) [CVPR24] 37.90 0.98 0.02
D-3DGS (Yang et al., 2024b) [CVPR24] 39.51 0.99 0.01
Table 5. Comparison of reconstruction methods on ZJU-MoCap (Peng et al., 2021b) (avatar), in terms of PSNR, SSIM, and LPIPS* (LPIPS $\times$ 1000). The results for non-GS methods are taken from (Lei et al., 2024a).
[…]
NeuralBody (Peng et al., 2021b) [CVPR21] 29.03 0.96 42.47
AnimNeRF (Peng et al., 2021a) [ICCV21] 29.77 0.96 46.89
PixelNeRF (Yu et al., 2021a) [ICCV21] 24.71 0.89 121.86
NHP (Kwon et al., 2021) [NeurIPS21] 28.25 0.95 64.77
HumanNeRF (Weng et al., 2022) [CVPR22] 30.66 0.97 33.38
Instant-NVR (Geng et al., 2023) [CVPR23] 31.01 0.97 38.45
GauHuman (Hu and Liu, 2024) [CVPR24] 31.34 0.97 30.51
3DGS-Avatar (Qian et al., 2024b) [CVPR24] 30.61 0.97 29.58
GART (Lei et al., 2024a) [CVPR24] 32.22 0.98 29.21
@@ 6.3. Performance Benchmarking: Dynamic Scenes
$\bullet$ Dataset: D-NeRF (Pumarola et al., 2021) dataset includes videos with 50 to 200 frames each, captured from unique viewpoints. It features synthetic, animated objects in complex scenes, with non-Lambertian materials. The dataset provides 50 to 200 training images and 20 test images per scene, designed for evaluating models in the monocular setting. The testing views are the same as the original paper (Pumarola et al., 2021).
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve five recent papers that model dynamic scenes with 3D GS (Yu et al., 2024b; Wu et al., 2024c; Lu et al., 2024a; Yang et al., 2024b, d), as well as six NeRF based approaches (Pumarola et al., 2021; Guo et al., 2023; Wang et al., 2023a).
$\bullet$ Evaluation Metric: The same metrics as in Sec. 6.2, i.e., PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018), are used for evaluation.
$\bullet$ Result: From Table 6.3 we can observe that 3D GS based methods outperform existing SOTAs by a clear margin. The static version of 3D GS (Kerbl et al., 2023) fails to reconstruct dynamic scenes, resulting in a sharp drop in performance. By modeling the dynamics, D-3DGS (Yang et al., 2024b) outperforms the SOTA method, FFDNeRF (Guo et al., 2023), by 6.83dB in terms of PSNR. These results indicate the effectiveness of introducing additional properties or structured information to model the deformation of Gaussians so as to model the scene dynamics.
6.4. Performance Benchmarking: Human Avatar
$\bullet$ Dataset: ZJU-MoCap (Peng et al., 2021b) is a prevalent benchmark in human modeling from videos, captured with 23 synchronized cameras at a $\times1024$ resolution. Six subjects (i.e., 377, 386, 387, 392, 393, and 394) are used for evaluation (Weng et al., 2022). The same testing views following (Geng et al., 2023) are adopted.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve three recent papers which model human avatar with 3D GS (Lei et al., 2024a; Hu and Liu, 2024; Qian et al., 2024b), as well as six human rendering approaches (Peng et al., 2021b; Weng et al., 2022; Geng et al., 2023).
$\bullet$ Evaluation Metric: PSNR, SSIM (Wang et al., 2004), and LPIPS* (Zhang et al., 2018) are used for measuring RGB rendering performance.
$\bullet$ Result: Table 6.3 presents the numerical results of top-leading solutions in human avatar modeling. We observe that introducing 3D GS into the framework leads to consistent performance improvements in both rendering quality and speed. For instance, GART (Lei et al., 2024a) outperforms current SOTA, Instant-NVR (Geng et al., 2023), by 1.21dB in terms of PSNR. Considering the enhanced fidelity, inference speed and editability, 3D GS based avatar modeling may revolutionize the field of 3D animation, interactive gaming, etc.
6.5. Performance Benchmarking: Surgical Scenes
3D reconstruction from endoscopic video is critical to robotic-assisted minimally invasive surgery, enabling preoperative planning, doctor training, etc.
$\bullet$ Dataset: EndoNeRF (Wang et al., 2022) dataset presents a specialized collection of stereo camera captures, comprising two samples of in-vivo prostatectomy. It is tailored to represent real-world surgical complexities, including challenging scenes with tool occlusion and pronounced non-rigid deformation. The same testing views as in (Zha et al., 2023) are used.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve three works which reconstruct dynamic 3D endoscopic scenes with GS (Huang et al., 2024b; Liu et al., 2024b; Zhao et al., 2024d), as well as three NeRF-based surgical reconstruction approaches (Wang et al., 2022; Zha et al., 2023).
$\bullet$ Evaluation Metric: PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) are adopted for evaluation.
Table 6. Comparison of reconstruction methods on EndoNeRF (Wang et al., 2022) (surgical scenes), in terms of PSNR, SSIM, and LPIPS. The results for non-GS methods are taken from (Liu et al., 2024b). FPS and GPU usage for training (Mem.) are measured using one 4090 GPU (Liu et al., 2024b).
@@ 6.5. Performance Benchmarking: Surgical Scenes
EndoNeRF (Wang et al., 2022) [MICCAI22] 36.06 0.93 0.09 0.04 19GB
EndoSurf (Zha et al., 2023) [MICCAI23] 36.53 0.95 0.07 0.04 17GB
LerPlane-9k (Yang et al., 2023b) [MICCAI23] 35.00 0.93 0.08 0.91 20GB
LerPlane-32k (Yang et al., 2023b) [MICCAI23] 37.38 0.95 0.05 0.87 20GB
Endo-4DGS (Huang et al., 2024b) [MICCAI24] 37.00 0.96 0.05 - 4GB
EndoGaussian (Liu et al., 2024b) [arXiv] 37.85 0.96 0.05 195.09 2GB
HFGS (Zhao et al., 2024d) [BMVC24] 38.14 0.97 0.03 - -
$\bullet$ Result: Table 6.5 shows that introducing the explicit representation of 3D Gaussians leads to several significant improvements. For instance, EndoGaussian (Liu et al., 2024b) outperforms a strong baseline, EndoSurf (Zha et al., 2023), among all metrics. In particular, EndoGaussian demonstrates an approximate 200-fold increase in speed while consumes just 10% of the GPU resources. These impressive results attest to the efficiency of GS-based methods, which not only expedite processing but also minimize GPU load, thus easing the demands on hardware. Such attributes are vitally significant for real-world surgical application deployment, where optimized resource usage can be a key determinant of practical utility.
7. Future Research Directions
@@
$\bullet$ Physics- and Semantics-aware Scene Representation. As a new, explicit scene representation technique, 3D Gaussian offers transformative potential beyond merely enhancing novel-view synthesis. It has the potential to pave the way for simultaneous advancements in scene reconstruction and understanding by devising physics- and semantics-aware 3D GS systems. While significant progress has been made in physics (Sec. 5.7) and semantics (Wang et al., 2024f; Qu et al., 2024; Ji et al., 2024d; Liao et al., 2024; Choi et al., 2024; Ji et al., 2024c) individually, there remains considerable untapped potential in their synergistic integration. This is poised to revolutionize a range of fields and downstream applications. For instance, incorporating prior knowledge such as the general shape of objects can reduce the need for extensive training viewpoints (Szymanowicz et al., 2024b; Charatan et al., 2024) while improving geometry/surface reconstruction (Guédon and Lepetit, 2024; Li et al., 2024f). A critical metric for assessing scene representation is the quality of its generated scenes, which encompasses challenges in geometry, texture, and lighting fidelity (Huang et al., 2024d; Chen et al., 2024a; Gao et al., 2023). By merging physical principles and semantic information within the 3D GS framework, one can expect that the quality will be enhanced, thereby facilitating dynamics modeling (Xie et al., 2024; Jiang et al., 2024c), editing (Cen et al., 2023; Zhou et al., 2024a), generation (Tang et al., 2024b; Yi et al., 2024), and beyond. In a nutshell, pursuing this advanced and versatile scene representation opens up new possibilities for innovation in computational creativity and practical applications across diverse domains.
@@
$\bullet$ Learning Physical Priors from Large-scale Data. As we explore the potential of physics- and semantics-aware scene representations, leveraging large-scale datasets to learn generalizable, physical priors emerges as a promising direction. The goal is to model the inherent physical properties and dynamics embedded within real-world data, transforming them into actionable insights that can be applied across various domains such as robotics and visual effects. Establishing a learning framework for extracting these generalizable priors enables the application of these insights to specific tasks in a few-shot manner. For instance, it allows for rapid adaptation to new objects and environments with minimal data input. Furthermore, integrating physical priors can enhance not only the accuracy and quality of generated scenes but also their interactive and dynamic qualities. This is particularly valuable in AR/VR environments, where users interact with virtual objects that behave in ways consistent with their real-world counterparts. However, the existing body of work on capturing and distilling physics-based knowledge from extensive 2D and 3D datasets remains sparse. Notable efforts in related area include the continuum mechanics based GS systems (Sec. 5.7), and the generalizable Gaussian representation based on multi-view stereo (Liu et al., 2024d). Further exploration on real2sim and sim2real might offer viable routes for advancements in this field.
@@
$\bullet$ Modeling Internal Structures of Objects with 3D GS. Despite the ability of 3D GS to produce highly photorealistic renderings, modeling internal structures of objects (e.g., for a scanned object in computed tomography) within the current GS framework presents a notable challenge. Due to the splatting and density control process, the current representation of 3D Gaussian is unorganized and cannot align well with the object’s actual internal structures. Moreover, there is a strong preference in various applications to depict objects as volumes (e.g., computed tomography). However, the disordered nature of 3D GS makes volume modeling particularly difficult. Li et al. (Li et al., 2023a) used 3D Gaussians with density control as the basis for the volumetric representation and did not involve the splatting process. X-Gaussian (Cai et al., 2024a) involves the splatting process for fast training and inference but cannot generate volumetric representation. Using 3D GS to model the internal structures of objects remains unanswered and deserves further exploration.
@@
$\bullet$ 3D GS for Simulation in Autonomous Driving and beyond. Collecting real-world datasets for autonomous driving is both expensive and logistically challenging, yet crucial for training effective image-based perception systems. To mitigate these issues, simulation emerges as a cost-effective alternative, enabling the generation of synthetic datasets across diverse environments. However, the development of simulators capable of producing photorealistic and diverse synthetic data is fraught with challenges. These include achieving a high level of quality, accommodating various control methods, and accurately simulating a range of lighting conditions. While early efforts (Zhou et al., 2024c; Yan et al., 2024a; Zhou et al., 2024e) in reconstructing urban/street scenes with 3D GS have been encouraging, they are just the tip of the iceberg in terms of the full capabilities. There remain numerous critical aspects to be explored, such as the integration of user-defined object models, the modeling of physics-aware scene changes (e.g., the rotation of vehicle wheels), and the enhancement of controllability and quality (e.g., in varying lighting conditions). Mastery of these capabilities would not only advance autonomous systems but also redefine computational understanding of physical spaces — a leap with implications for world models, spatial intelligence, embodied AI, and beyond.
@@
$\bullet$ Empowering 3D GS with More Possibilities. Despite the significant potential of 3D GS, the full scope of applications for 3D GS remains largely untapped. A promising avenue for exploration involves augmenting 3D Gaussians with additional attributes (e.g., linguistic and spatiotemporal properties as mentioned in Sec. 4.5) and introducing structured information (e.g., spatial MLPs and grids as mentioned in Sec. 4.6), tailored for specific applications. Moreover, recent studies have begun to unveil the capability of 3D GS in several domains, e.g., point cloud registration (Chang et al., 2024), image representation and compression (Zhang et al., 2024d), and fluid synthesis (Feng et al., 2024a). These findings highlight a significant opportunity for interdisciplinary scholars to explore 3D GS further.
8. Conclusions